# Logical perspectives on learning statistical objects

### v1

@@ Abstract
We consider the relationship between learnability of a “base class” of functions on a set $X$, and learnability of a class of statistical functions derived from the base class. For example, we refine results showing that learnability of a family $h_{p}:p\in Y$ of functions implies learnability of the family of functions $h_{\mu}=\lambda p:Y.E_{\mu}(h_{p})$, where $E_{\mu}$ is the expectation with respect to $\mu$, and $\mu$ ranges over probability distributions on $X$. We will look at both Probably Approximately Correct (PAC) learning, where example inputs and outputs are chosen at random, and online learning, where the examples are chosen adversarily. We establish improved bounds on the sample complexity of learning for statistical classes, stated in terms of combinatorial dimensions of the base class. We do this by adapting techniques introduced in model theory for “randomizing a structure”. We give particular attention to classes derived from logical formulas, and relate learnability of the statistical classes to properties of the formula. Finally, we provide bounds on the complexity of learning the statistical classes built on top of a logic-based hypothesis class.
@@ 1 Introduction
We analyze these two statistical classes by embedding in the randomization of a hypothesis class, inspired by work in model theory (Keisler, 1999; Ben Yaacov & Keisler, 2009; Ben Yaacov, 2009). In those works, we move from a base structure to another structure, its randomization. In the PAC learning scenario, we can apply prior work in model theory (Ben Yaacov, 2009) to conclude that when the base class is PAC learnable, the randomization class is PAC learnable. We show that this implies preservation of learnability of both distribution classes. We refine these arguments in several directions: to apply to new statistical classes, to deal with real-valued functions in the base class, and to get bounds on the number of samples needed to learn. We also examine whether the same phenomenon applies to other learning scenarios: e.g. realizable PAC learning, online learning.
We will pay particular attention to the setting where classes come from varying parameters within a logical formula over an infinite structure: definable families. Standard examples of learnable classes – e.g. rectangles, families of regions defined by polynomials of fixed degree – fit into this framework. We show that for many common structures, all the statistical classes built on top of definable families are learnable.
Contributions: preservation, sample complexity, and decidability. Our first results concern whether learnability is preserved when moving from a base hypothesis class to the corresponding statistical class. For agnostic learning, we provide positive results on preservation, accompanied by sample complexity bounds for the statistical class, stated in terms of combinatorial dimensions of the base class. For realizable learning, we show negative results. A high-level overview is in Table 1. The formal definitions are in the next sections.
@@
Preserved (Cor. 2) Not Preserved for Dual Dist. Class (Prop. 7)
PAC Finite FatSh Bartlett & Long (1995) Finite OIG dimension Attias et al. (2023)
Preserved (Thm 5) Not Preserved for Distr. or Dual Distr. Class (Prop. 5)
@@
The results above concern sample complexity – the number of samples needed to learn a statistical object within a given tolerance and a given confidence. We also examine computational complexity of learning, focusing exclusively on classes defined by first-order logic formulas. We show that under some decidability-related hypothesis on the logical structure, we can effectively estimate the value of a statistical function on a new input point, within a given tolerance.
Organization. Section 2 reviews different flavors of learnability that we study here. Section 3 reviews how learnability and hypothesis classes arise from logic. Section 4 defines the notion of “statistical class built over a base class” that will be our central object of study.
Section 5 studies preservation of PAC learnability for statistical classes, in two variations of the learning set up: agnostic and realizable. Section 6 performs the same study for online learning. The bounds in the bulk of this work are on the sample complexity of learning: how many samples do we need to make a prediction with a certain confidence. Section 7 briefly discusses the computational complexity of learning. We discuss related work in Section 8. We close with conclusions in Section 9.
We defer the proofs of a few propositions and lemmas to the appendix.
@@ 2 Preliminaries
Functions in a hypothesis class ${\mathcal{H}}$ are expressed as $h_{p}$ where $p$ ranges over the parameter space of the class $Y$. A family of functions on $X$ indexed by a set $Y$ can be considered as a single function from $X\times Y$ to $[0,1]$. This corresponds naturally to a function $Y\times X\to[0,1]$, and thus also defines a class of functions on $Y$ indexed by $X$, the dual class of ${\mathcal{H}}$.
PAC learning. We recall the standard notion of a function class being learnable (Kearns & Schapire, 1994) from random supervision: Probably Approximately Correct or PAC learnable below. Fixing $X$, let $Z=X\times[0,1]$. We call the elements of $Z$ samples. For each hypothesis $h\in{\mathcal{H}}$ and sample $z=(x,y)$, let $l_{h}(z)=(h(x)-y)^{2}$: this is the loss of using this hypothesis at sample $z$. For a distribution $P$ over $Z$, we let $\mathrm{ExpLoss}_{P}(h)$ be the expected loss of $h$ with respect to $P$. We let $\mathrm{BestExpLoss}(P)$ be the infimum of $\mathrm{ExpLoss}_{P}(h)$ over every $h\in{\mathcal{H}}$.
@@
$P^{n}(\vec{z}~{}|~{}\mathrm{ExpLoss}_{P}(A(\vec{z}))\leq\mathrm{BestExpLoss}_{P}+\epsilon)\geq 1-\delta.$
@@
Fact 1 (See, e.g. Shalev-Shwartz & Ben-David (2014)).
@@
Dimensions of real-valued families. Each of our notions of learnability corresponds to a combinatorial dimension of the class. For agnostic PAC learning, the characterization involves the fat-shattering definition originating in (Kearns & Schapire, 1994).
Definition 1 (Fat shattering).
@@
Fact 4 (Bartlett & Long (1995, Thm 14)).
@@
Definition 2 (Sequential fat-shattering dimension and Littlestone dimension).
@@
Branches $B\in\{-1,1\}^{d}$ can also be viewed as functions $B:\{0,\dots,d-1\}\to\{-1,1\}$, with $B(t)$ the $t$th entry of $B$.
@@
Fact 5 (Alon et al. (2021, Theorem 12.1)).
@@
Fact 6 (Rakhlin et al. (2015b, Part of Proposition 9)).
@@
Fact 7.
Kleer & Simon (2023) A function class is agnostic PAC learnable exactly when its dual class is.
@@
3 Learnable classes from logic
We will now review a way to get hypothesis classes using logic, by looking at the hypothesis class associated to a formula and a partition of its free variables.
We deal with the standard definition of first-order logic, parameterized by a vocabulary $L$, consisting of relation and function symbols of different arities, including zero (e.g. for constant symbols). The first-order terms are built up from variables and constant symbols by applying function symbols: e.g $f(g(c),v)$. The first-order logic atomic formulas are of the form $R(\tau_{1}\ldots\tau_{n})$ where $R$ has arity $n$ and $\tau_{1}\ldots\tau_{n}$ are terms, along with $\tau_{1}=\tau_{2}$ where $\tau_{i}$ are terms. The first-order logic formulas are built up from atomic formulas via Boolean operations ,,$\wedge$,$\vee$,$\neg$ , ∨ , ¬ along with quantifications $\exists x~{}\phi$, $\forall x~{}\phi$. The semantics are standard (Hodges, 1993).
A structure $\mathfrak{M}$ for language $L$ comes with a domain or universe, which we denote as $M$, along with an interpretation of each symbol in $L$: for example, a unary predicate will be interpreted by a subset of $M$. We fix a structure $\mathfrak{M}$ with vocabulary $L$ and domain $M$. We consider any first-order logic formula $\phi(x_{1}\ldots x_{j};y_{1}\ldots y_{k})$ where in the notation we also fix a partition of the free variables into $\vec{x}$ and $\vec{y}$.
Definition 3 (Concept class of a (variable-partitioned) first-order formula).
A partitioned formula $\phi$ defines in the obvious way a function ${\mathcal{H}}_{\phi}$ on $M^{j}\times M^{k}$ to $\{0,1\}$, which can be considered as a family of functions on range space $M^{j}$ indexed by parameters in $M^{k}$. By convention, we write the variables that are considered parameters second in a partition. Thus the canonical concept class associated with $\phi(x_{1}\ldots x_{j};y_{1}\ldots y_{k})$, denoted ${\mathcal{C}}_{\phi}$, has $y_{1}\ldots y_{k}$ as parameters, with each parameter value giving a concept over $M^{j}$.
Definition 4 (NIP partitioned formulas and structures).
Fixing $\mathfrak{M}$ we say that a partitioned first-order logic formula $\phi(\vec{x},\vec{y})$ in the language of $\mathfrak{M}$ is NIP if the corresponding family ${\mathcal{C}}_{\phi}$ has finite VC dimension, and say that $\mathfrak{M}$ is NIP if every partitioned first-order formula has this property.
NIP stands for “Not the Independence Property”, referring to an independent family of sets, one where every Boolean combination of family members is non-empty. So NIP partitioned formulas are those where the corresponding family does not have arbitrarily large independent subfamilies.
When we have established that a structure is NIP, we know how to use any partitioned formula to get a PAC learnable class. There are many tools available to verify that a structure is NIP, and examples include (Simon, 2015):
  • The real ordered group $({\mathcal{R}},+,<)$ and the real ordered field $({\mathcal{R}},+,\cdot,<)$
  • The real exponential field $({\mathcal{R}},+,\cdot,\lambda x.e^{x},<)$
  • The complex field $(\mathbb{C},+,\cdot)$
  • The infinite complete binary tree with predicates for the two successor relations.
  • Integer additive arithmetic $(\mathbb{N},+,<)$
Thus any definable family in each of these structures is learnable. Logic-based concept classes include many of the familiar examples of finite VC-dimension, such as rectangles. Concrete bounds on the corresponding dimensions, and thus the sample complexity of PAC learning, have been established in many cases (Karpinski & Macintyre, 1997; Chistikov et al., 2022).
3.1 Real-valued functions from continuous logic
We now discuss how a structure can provide a collection of learnable real-valued functions. For this we utilize continuous logic (CL for short), defined over structures where the atomic predicates are real-valued: that is, a $k$-ary real-valued predicate is just a function from $k$-tuples in the model into the reals. We may also have $k$-ary function symbols $f(x_{1}\ldots x_{k})$, which are associated to functions on elements in the model, as in first-order logic. We write $\mathfrak{C}$ for such a structure.
We will be able to make do with a simplified presentation of the syntax and semantics of CL, since many aspects in the logic literature – e.g. metrics, variations of the compactness theorem, Lowenheim-Skolem theorem, ultraproducts – will not be needed for the results we state. We form the set of basic formulas over $\mathfrak{C}$, following (Hart, 2023). For any valuation of the free variables, the semantics will associate these formulas with real numbers.
  • Terms are built up from variables using function symbols as in first-order logic. For any $j$-ary predicate $P$ in the signature and terms $\tau_{1}\ldots\tau_{j}$ we have a basic formula $P(\tau_{1}\ldots\tau_{j})$. The semantics under a valuation for the variables just evaluates each $\tau_{i}$ to get a vector of elements in the model, and then applies the real-valued function associated to $P$.
  • If $f:[0,1]^{n}\rightarrow[0,1]$ is continuous, and $\phi_{1}\ldots\phi_{n}$ are basic formulas, then $f(\phi_{1},\ldots\phi_{n})$ is a basic formula, with the obvious semantics.
  • If $\phi(x,\vec{y})$ is a basic formula with $x$ a variable, then $\sup_{x}\phi$ and $\inf_{x}\phi$ are basic formulas.
Notice that if we start with a classical first-order structure, the range of each basic formula is a finite set, and we are not really gaining expressiveness over first-order logic: a formula will just partition tuples in the model into first-order definable sets and then associate each partition with a real number.
A formula of continuous logic extends the basic formulas above by closing under an infinite summation construct: if we have formulas $\phi_{i}:i\in\mathbb{N}$, then we can form a new formula $\Sigma_{i}\frac{\phi_{i}}{2^{i}}$. The semantics are again obvious: inductively, we can show that the range of each formula is bounded, which guarantees that the sum converges.
Definition 5 (Hypothesis class of a real-valued formula).
Consider a CL formula $\phi(x_{1}\ldots x_{j};y_{1}\ldots y_{k})$ with a partition of the free variables into $\vec{x}$ and $\vec{y}$. From this we get a function from $M^{j}\times M^{k}\rightarrow[0,1]$, which we can consider as a hypothesis class ${\mathcal{H}}_{\phi}$ over $M^{j}$ indexed by $M^{k}$.
We say that a CL structure $\mathfrak{C}$ is NIP if for every partitioned real valued formula $\phi(\vec{x};\vec{y})$, the corresponding family of real valued functions (as we vary $\vec{y}$ over the structure), has bounded fat-shattering dimension – or equivalently, by Fact 4, the family is agnostic PAC learnable.
Any first-order structure can be viewed as a continuous logic structure: we have $\{0,1\}$ valued basic predicates, and we can also consider equality as a basic predicate, but we have many more formulas. We write $\mathfrak{C}(\mathfrak{M})$ for this continuous logic structure. It is easy to prove that if $\mathfrak{C}$ is NIP, then so is $\mathfrak{C}(\mathfrak{M})$. That is, throwing in these additional formulas does not give us function classes that are not agnostic PAC learnable: see Appendix B.
Proposition 2.
If a first-order structure $\mathfrak{M}$ is NIP, then for any general continuous logic partitioned formula $\phi$, ${\mathcal{H}}_{\phi}$ has finite fat-shattering dimension and is thus agnostic PAC learnable. Thus $\mathfrak{C}(\mathfrak{M})$ is also NIP.
Proposition 2 gives a way of using first-order structures to get real-valued hypothesis classes that are well-behaved in terms of learning. It is particularly useful when applying continuous logic to first-order structures over the reals. Let $\mathfrak{M}$ be a classical first-order structure on ${\mathcal{R}}$ in a signature including the ordering $<$.
Lemma 1.
Let $f:{\mathcal{R}}^{n}\to[0,1]$ be a definable function in $\mathfrak{M}$: one whose graph is defined by a formula $\phi(x_{1}\ldots x_{n},y)$. Then there is a formula in continuous logic over $\mathfrak{C}(\mathfrak{M})$ that has $f$ as its semantic function.
See Appendix C for the simple proof. Combining the previous proposition and lemma, we can get many NIP CL-definable function classes since there are multiple NIP structures over the reals (van den Dries, 1998).
Example 3.1. Consider the hypothesis class ${\mathcal{H}}_{\vec{p}}$ consisting of rational functions $\frac{P(x)}{Q(x)}$, where $P(x)$ and $Q(x)$ are real polynomials restricted to an interval where $Q$ has no zeros (with the function defined to be zero off this interval), the degrees of $P$ and $Q$ are at most $5$, and the value of the quotient function on the interval is in $[0,1]$. This is a class definable over the real field by a formula $\phi(x,y;\vec{p})$ where $\vec{p}$ are parameters representing the coefficients of $P$ and $Q$, $x$ represents the function input and $y$ the output. Thus since the real field is known to be an NIP expansion of $({\mathcal{R}},<)$ (van den Dries, 1998), Lemma 1 implies that this class of real-valued functions indexed by $\vec{p}$ is agnostic PAC learnable.
We emphasize that every concept class can be trivially seen as the concept class associated with a logical formula in a structure. We just take a two-sorted structure, with one sort being the parameter space of the concept, the other being the range space. Similarly, every real-valued hypothesis class can be trivially seen as being given by some continuous logic formula. We do not formalize these comments here, since most of our proofs can be translated directly from logic-based classes to general ones.
The logical perspective allows us to generate many families of learnable examples. It also gives an elegant way of seeing duality: if our hypothesis class is given by a partitioned formula $\phi(\vec{x};\vec{y})$, then moving to the dual just swaps the role of $\vec{y}$ and $\vec{x}$.
@@ 4 Hypothesis classes from statistical objects
Definition 6 (Distribution class and Dual Distribution class).
Given a concept class ${\mathcal{C}}$ on $X$, parameterized by $Y$ the distribution function class of ${\mathcal{C}}$, denoted ${\mathcal{D}istr}_{{\mathcal{C}}}$, is a real-valued hypothesis class on the same range space $X$, consisting of the hypotheses $h_{\mu}$ indexed by the distributions $\mu$ on $Y$ such that for each $x$, the set of $p\in Y$ such that $h_{p}(x)=1$ is measurable. The hypothesis $h_{\mu}$ is defined as the function mapping $x$ to the probability, with respect to $\mu$, that $h_{p}(x)=1$.
@@
We extend these definitions to work on top of a real-valued function class ${\mathcal{H}}$, replacing probability under $\mu$ with expectation. For example, for the distribution function class, the functions are indexed by distributions $\mu$ on $Y$ again, but now we restrict to distributions $\mu$ such that each $h\in{\mathcal{H}}$ is a $\mu$ measurable function, and $h_{\mu}$ maps $h\in{\mathcal{H}}$ to $E_{\mu}(h)$.
@@
The dual distribution class of ${\mathcal{H}}$ will consist of “random elements of the plane”: functions parameterized by distributions $\nu$ over the plane. Each such $\nu$ induces a function $h_{\nu}$ on rectangles, mapping each rectangle to its $\nu$-probability. We can equivalently consider $h_{\nu}$ as a real-valued function on $4$-tuples. In learning such a function $h_{\nu}$, we will be given supervision in the form of a sequence of pairs, each pair consisting of a rectangle (or $4$-tuple) and the $\nu$-probability that a point is in the rectangle.
In contrast, the distribution class of ${\mathcal{H}}$ will consist of “random rectangles”: functions parameterized by distributions $\mu$ over $4$-tuples. Given such a $\mu$, we have a function $h_{\nu}$ that maps a real pair $\vec{x}$ to the $\nu$ probability that $\vec{x}$ is in rectangle $h_{\mu}$. In learning such a function, our supervision will consist of a point in the real plane and the probability that the point is in random rectangle $\mu$.

### v2

# From learnable objects to learnable random objects

@@ Abstract
We consider the relationship between learnability of a “base class” of functions on a set $X$, and learnability of a class of statistical functions derived from the base class. For example, we refine results showing that learnability of a family $h_{p}:p\in Y$ of functions implies learnability of the family of functions $h_{\mu}=\lambda p:Y.E_{\mu}(h_{p})$, where $E_{\mu}$ is the expectation with respect to $\mu$, and $\mu$ ranges over probability distributions on $X$. We will look at both Probably Approximately Correct (PAC) learning, where example inputs and outputs are chosen at random, and online learning, where the examples are chosen adversarily. For agnostic learning, we establish improved bounds on the sample complexity of learning for statistical classes, stated in terms of combinatorial dimensions of the base class. We connect these problems to techniques introduced in model theory for “randomizing a structure”. We also provide counterexamples for realizable learning, in both the PAC and online settings.
@@ 1 Introduction
We analyze these two statistical classes using a broader result about random hypotheses classes inspired by work in model theory (Keisler, 1999; Ben Yaacov and Keisler, 2009; Ben Yaacov, 2009). In the PAC learning scenario, we can apply prior work in model theory (Ben Yaacov, 2009) to conclude that when a random family of hypothesis classes is uniformly learnable, then the expectation of this class is also learnable. We show that this implies preservation of learnability of both distribution classes. We refine these arguments in several directions: to apply to new statistical classes, to deal with real-valued functions in the base class, and to get bounds on the number of samples needed to learn. We also examine whether the same phenomenon applies to other learning scenarios: e.g. realizable PAC learning, online learning.
Contributions: preservation and sample complexity. Our results concern whether various notions of learnability are preserved when moving from a base hypothesis class to the corresponding statistical class. For agnostic learning, both PAC and online, we provide positive results on preservation, accompanied by sample complexity bounds for the statistical class, stated in terms of combinatorial dimensions of the base class. For realizable learning, we show negative results. A high-level overview is in Table 1, which highlights the positive and negative preservation results we prove in moving from a base class to a statistical class, and the combinatorial dimension that characterizes learnability. In each of the positive results, we supply sample complexity bounds, not listed in the table. The formal definitions are in the next sections.
@@
Preserved (Cor. 35) Not Preserved for Dual Dist. Class (Prop. 51)
PAC Finite FatSh Bartlett and Long (1995) Finite OIG dimension Attias et al. (2023)
Preserved (Thm 15) Not Preserved for Distr
or Dual Distr. Class (Prop. 30)
@@
Organization. Section 2 reviews different flavors of learnability that we study here. Section 3 defines the notion of “statistical class built over a base class” that will be our central object of study. Section 4 studies preservation of PAC learnability for statistical classes, in two variations of the learning set up: agnostic and realizable. Section 5 performs the same study for online learning. We discuss related work in Section 6. We close with conclusions in Section 7. We defer the proofs of a few propositions and lemmas to the appendix.
@@ 2 Preliminaries
Functions in a hypothesis class ${\mathcal{H}}$ are expressed as $h_{p}$ where $p$ ranges over the parameter space of the class $Y$. We write ${\mathcal{C}}_{p}$ for a concept that is in a concept class ${\mathcal{C}}$.
A family of functions on $X$ indexed by a set $Y$ can be considered as a single function from $X\times Y$ to $[0,1]$. This corresponds naturally to a function $Y\times X\to[0,1]$, and thus also defines a class of functions on $Y$ indexed by $X$, the dual class of ${\mathcal{H}}$.
PAC learning. We recall the standard notion of a function class being learnable (Kearns and Schapire, 1994) from random supervision: Probably Approximately Correct or PAC learnable below. Fixing $X$, let $Z=X\times[0,1]$. We call the elements of $Z$ samples. For each hypothesis $h\in{\mathcal{H}}$ and sample $z=(x,y)$, let $l_{h}(z)=(h(x)-y)^{2}$: this is the loss of using this hypothesis at sample $z$. For a distribution $P$ over $Z$, we let $\mathrm{ExpLoss}_{P}(h)$ be the expected loss of $h$ with respect to $P$. We let $\mathrm{BestExpLoss}(P)$ be the infimum of $\mathrm{ExpLoss}_{P}(h)$ over every $h\in{\mathcal{H}}$.
@@
$P^{n}(\vec{z}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathrm{ExpLoss}_{P}(A(\vec{z}))\leq\mathrm{BestExpLoss}_{P}+\epsilon)\geq 1-\delta.$
@@
Fact 1 (See, e.g. Shalev-Shwartz and Ben-David (2014))
@@
Dimensions of real-valued families. Each of our notions of learnability corresponds to a combinatorial dimension of the class. For agnostic PAC learning, the characterization involves the fat-shattering definition originating in (Kearns and Schapire, 1994).
Definition 1 (Fat shattering)
@@
Fact 4 (Bartlett and Long (1995, Thm 14))
@@
Definition 2 (Sequential fat-shattering dimension and Littlestone dimension)
@@
Branches $B\in\{-1,1\}^{d}$ can also be viewed as functions $B:\{0,\dots,d-1\}\to\{-1,1\}$, with $B(t)$ the $t^{th}$ entry of $B$.
@@
Fact 5 (Alon et al. (2021, Theorem 12.1))
@@
Fact 6 (Rakhlin et al. (2015b, Part of Proposition 9))
@@
Fact 7
Kleer and Simon (2023) A function class is agnostic PAC learnable exactly when its dual class is.
@@ 3 Hypothesis classes from statistical objects
Definition 4 (Distribution class and Dual Distribution class)
Given a concept class ${\mathcal{C}}$ on $X$, parameterized by $Y$ the distribution function class of ${\mathcal{C}}$, denoted ${\mathcal{D}istr}_{{\mathcal{C}}}$, is a real-valued hypothesis class on the same range space $X$, consisting of the hypotheses $h_{\mu}$ indexed by the distributions $\mu$, with $\sigma$-algebra $\Sigma_{\mu}$, on $Y$ such that for each $x$, the set of $p\in Y$ such that $c_{p}(x)=1$ is measurable.
The hypothesis $h_{\mu}$ is defined as the function mapping $x$ to the probability, with respect to $\mu$, that $c_{p}(x)=1$. Note that, WLOG, we can assume that the $\Sigma_{\mu}$ is the $\Sigma$-algebra generated by $\{\{p\in Y:c_{p}(x)=1\}:x\in X\}$, since $h_{\mu}$ is completely determined by the restriction of $\mu$ to this $\Sigma$-algebra.
@@
We extend these definitions to work on top of a real-valued function class ${\mathcal{H}}$, replacing probability under $\mu$ with expectation. For the distribution function class, the functions are indexed by distributions $\mu$ on $Y$ again, but now we restrict to distributions $\mu$ such that each $h\in{\mathcal{H}}$ is a $\mu$ measurable function, and $h_{\mu}$ maps $h\in{\mathcal{H}}$ to $E_{\mu}(h)$. As with the concept class, we can WLOG assume that $\Sigma_{\mu}$ is the $\sigma$-algebra generated by the sets $\{y:h_{y}(x)\in I\}$ where $I$ is an interval, $x\in X$.
@@
The dual distribution class of ${\mathcal{H}}$ will be parameterized by “random elements of the plane”: functions induced by distributions $\mu$ over the plane. Each such $\mu$ induces a function $h_{\nu}$ on rectangles, mapping each rectangle to its $\nu$-probability. We can equivalently consider $h_{\mu}$ as a real-valued function on $4$-tuples. In learning such a function $h_{\mu}$, we will be given supervision in the form of a sequence of pairs, each pair consisting of a rectangle (or $4$-tuple) and the $\mu$-probability that a point is in the rectangle.
In contrast, the distribution class of ${\mathcal{H}}$ is parameterized by “random rectangles”: functions induced by distributions $\nu$ over $4$-tuples. Given such a $\nu$, we have a function $h_{\nu}$ that maps a real pair $\vec{x}$ to the $\nu$ probability that $\vec{x}$ is in rectangle $h_{\nu}$. In learning such a function, our supervision will consist of a point in the real plane and the probability that the point is in random rectangle $\nu$.
Example 3.2. Consider the hypothesis class ${\mathcal{H}}$ consisting of rational functions $\frac{P(x)}{Q(x)}$, where $P(x)$ and $Q(x)$ are real polynomials restricted to an interval where $Q$ has no zeros (with the function defined to be zero off this interval), the degrees of $P$ and $Q$ are at most $5$, and the value of the quotient function on the interval is in $[0,1]$. This is a class with parameters $\vec{p}$ representing the coefficients of $P$ and $Q$. We can easily show (see the appendix for a broader argument) that this class of real-valued functions indexed by $\vec{p}$ is agnostic PAC learnable.
The dual distribution class is parameterized by “random arguments” to these functions: with each distribution inducing a function on parameters given by integrating against the distribution. The distribution class will be parameterized by “random coefficients”: each distribution will induce a function on arguments $x$ obtained by integrating the random function against the distribution on parameters.
@@
The motivation in Hu et al. (2022) was from databases: we want to learn the selectivity of a table in a databases. Given some examples of the density of the table in some intervals, we want to predict the density in a new interval. Hu et al. (2022) deals with computational complexity as well as sample complexity, but here we will only be interested in sample complexity bounds.
3.1 Expectation of a Measurable Family of Hypothesis Classes